# A Watermark for Low-entropy and Unbiased Generation in Large Language Models

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have highlighted the risk of misusing them, raising the need for accurate detection of LLM-generated content. In response, a viable solution is to inject imperceptible identifiers into LLMs, known as watermarks. Previous work demonstrates that unbiased watermarks ensure unforgeability and preserve text quality by maintaining the expectation of the LLM output probability distribution. However, previous unbiased watermarking methods suffer from one or more of the following issues: (1) requiring access to white-box LLMs during detection, (2) incurring long detection time, (3) being not robust against simple watermarking attacks, (4) failing to provide statistical guarantees for the type II error of watermark detection, and (5) being not statistically unbiased for low-entropy scenarios, which hinder their deployment in practice. This study proposes the Sampling One Then Accepting (STA-1) method, a watermark that can address all of these issues. Moreover, we discuss the tradeoff between watermark strength and text quality for unbiased watermarks. We show that in low-entropy scenarios, unbiased watermarks face a tradeoff between watermark strength and the risk of unsatisfactory outputs. Experimental results on both low-entropy and high-entropy datasets demonstrate that STA-1 achieves text quality and watermark strength comparable to existing unbiased watermarks, with a low risk of unsatisfactory outputs. Implementation codes for this study are available online. [^1] https://github.com/djwei96/STA
@@ 1 Introduction
Large language models (LLMs) are large-scale deep learning models that can understand and generate natural languages by learning from a large amount of textual data. Typical generative LLMs, such as ChatGPT (Ouyang et al., 2022) and LLaMA (Touvron et al., 2023), can answer questions, translate languages, and create codes with qualities comparable to humans. As LLMs can generate contents more efficiently at a lower cost compared to humans, the risk of LLMs being employed to generate biased, fake, or malicious contents is also increasing (Mirsky et al., 2023). For example, LLMs may exhibit biased information against underrepresented groups of people (Abid et al., 2021; Fang et al., 2024), create misinformation (Pan et al., 2023), and harm academic integrity (Zhao et al., 2023). To reduce the harm caused by LLMs, identifying LLM-generated content precisely and efficiently becomes a crucial issue (Kirchenbauer et al., 2023a).
Watermarks are identifiers imperceptible to humans but detectable by certain models (Liu et al., 2023b). In the era of LLMs, a watermarking method is a strategy to control the randomness of token generation by LLMs (Kirchenbauer et al., 2023a; Fernandez et al., 2023), with the randomness preserved confidentially by LLM owners. In practice, watermarks should have unforgeability against deciphering watermarking generation attacks (Liu et al., 2023b). A watermarking method demonstrates the ability against forgeries if it can hide the distinguishability between the original unwatermarked text and its watermarked counterpart (Christ et al., 2023; Liu et al., 2023b). Thus, it is required that a watermarking method adjusts the probability distribution while maintaining the same expectation as the unwatermarked distribution (Hu et al., 2024; Kuditipudi et al., 2023), defined as unbiased watermarks.
We conclude the research gaps in Table 5 in Appendix A. In response to the challenges, we propose the Sampling One Then Accepting (STA-1) method that can simultaneously overcome the above gaps. STA-1 traces back to the original watermarking method (denoted as KGW) (Kirchenbauer et al., 2023a) where the token set is divided into a green and a red list at each generation step. Instead of raising logits in the green list, STA-1 samples a token from the original probability distribution and accepts it if it is in the green list. If the sampled token is in the red list, it resamples another token and accept it. By counting the number of green list tokens, it employs the $z$-test for watermark detection, which naturally addresses the issues encountered with the LLR test and eliminates the need for prompts and white-box LLMs in detection. Simultaneously, the detection of STA-1 is efficient and only requires $O(m)$ time complexity, where $m$ is the number of tokens. The STA-1 method is also robust against simple insertion and removal attacks because changing tokens can only affect the detection score around these tokens. Meanwhile, we prove that STA-1 is unbiased and provides statistical guarantees for the type II error. More interestingly, the bounds are linked to the Gini index of the probability distribution, which is a common metric in machine learning (Breiman, 2017) compared to the proposed Spike entropy in previous work (Kirchenbauer et al., 2023a).
In this study, we also clarify the watermark strength and text quality tradeoff in unbiased watermarks. The KGW method faces a tradeoff between watermark strength and text quality, which means a higher detection power results in a lower text quality (Kirchenbauer et al., 2023a). Previous work claims that unbiased watermarks can avoid this tradeoff given the preserved text quality by maintaining the expectation of probability distribution (Hu et al., 2024). We challenge this claim by considering a simple low-entropy scenario, where we show that unbiased watermarks still face a tradeoff between watermark strength and text quality. However, under the same expectation constraint, the text quality is related to the risk of unsatisfactory outputs. Specifically, unsatisfactory outputs in low-entropy scenarios represent that the watermarking method alters probability distribution too much such that high-probability tokens cannot be sampled at risk. We discuss the risk via the variance of the probability after altering, which is a common practice of risk-return analysis (Sharpe, 1998). We prove that STA-1 is less risky than previous unbiased watermarks. Moreover, we propose STA-M, an extension of STA-1, by setting up a threshold for entropy in generation (Lee et al., 2023; Wang et al., 2023) and sampling more times for high-entropy steps. Although STA-M is not unbiased theoretically, it allows higher watermark strength with small performance shifts empirically. Also, STA-M is robust against various watermarking attacks. We summarize our contributions as follows:
@@
1. We propose STA-1, an unbiased watermarking method that is practical and has statistical guarantees on type II error of watermark detection. Moreover, we introduce STA-M, an extension of STA-1 that enhances watermark strength with low text quality shifts.
2. We clarify the watermark strength and text quality tradeoff in unbiased watermarks. In low-entropy scenarios, the text quality is related to the risk of unsatisfactory outputs. We show that STA-1 has a lower risk theoretically compared to other unbiased watermarks.
@@
3. Experimental results on public low-entropy and high-entropy datasets empirically show that STA-1 achieves comparable performances against other unbiased watermarks and has a low risk of unsatisfactory outputs. Meanwhile, STA-M demonstrates high watermark strength in the low-entropy dataset and robustness against different watermarking attacks.
@@ 2 Preliminary
Notations. We follow notations in previous work (Kirchenbauer et al., 2023a; Hu et al., 2024) to represent the generation task of LLMs. Let $P_{M}$ denote a pretrained LLM and $\mathcal{V}$ is the overall token (vocabulary) set. An example token set contains more than 50,000 tokens ($|\mathcal{V}|>50000$) (Radford et al., 2019). For simplicity, we use Python-style notation for an ordered token sequence, where $x^{-m:n}=(x^{-m},x^{-m+1},\cdots,x^{n})$, $m$ and $n$ are integers. In a typical LLM generation task, an LLM receives a sequence of $N_{p}+1$ tokens $x^{-N_{p}:0}$, known as a prompt, and outputs a sequence of $T$ tokens $x^{1:T}$ step by step. At step $t$, the probability of each token in the token set $\mathcal{V}$ is given by the conditional distribution $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$. The LLM generation follows an autoregressive fashion, where the joint probability of the generated tokens are as $P_{M}(x^{1:T}|x^{-N_{p}:0})=\prod_{t=1}^{T}P_{M}(x^{t}|x^{-N_{p}:(t-1)})$.
When applying watermarking techniques, the LLM employs a private key $k$ to adjust the conditional distribution from $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ to $P_{M,w}(x^{t}|x^{-N_{p}:(t-1)};k)$, where $P_{M,w}$ indicates a watermarked model and the private key $k$ is randomly selected from a key space $K$ according to a known distribution $P_{K}(k)$. An unbiased watermark requires that the expectation of the watermarked distribution equals that of the original distribution (Hu et al., 2024), defined as follows.
@@
$\mathbb{E}_{k\sim P_{K}(k)}\left[P_{M,w}(x^{t}|x^{-N_{p}:(t-1)};k)\right]=P_{M}(x^{t}|x^{-N_{p}:(t-1)}),$ (1)
@@
Previous distribution reweighting methods. Since controlled sampling can be viewed as a special case of distribution reweighting, we build our analysis framework based on distribution reweighting. Formally, a reweighting function $R_{k}:\mathcal{P}_{\mathcal{V}}\rightarrow\mathcal{P}_{\mathcal{V}}$ maps from $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ to $P_{M,w}(x^{t}|x^{-N_{p}:(t-1)};k)$, where $\mathcal{P}_{\mathcal{V}}$ denotes the probability distribution space over the vocabulary set $\mathcal{V}$. A reweighting method $R:K\times\mathcal{P}_{\mathcal{V}}\rightarrow\mathcal{P}_{\mathcal{V}}$ contains all realized reweighting functions $R_{k}$ among the key space $k\in K$. Following Definition 1, $R$ is an unbiased reweighting method if $\mathbb{E}_{k\sim P_{K}(k)}\left[R_{k}(P_{M})\right]=P_{M}$. Next, we introduce previous watermarking methods (Kirchenbauer et al., 2023a; Hu et al., 2024; Wu et al., 2024; Kuditipudi et al., 2023) and refer readers to Appendix B for more details.
KGW (Kirchenbauer et al., 2023a): The KGW method (Kirchenbauer et al., 2023a) randomly splits the vocabulary set $\mathcal{V}$ into a green list and a red list based on a uniformly distributed key $k$. The soft KGW method adds a predefined constant $\delta$ to the green list tokens’ logits while keeping the red list tokens’ logits fixed.
Dipmark (Wu et al., 2024) and $\gamma$-reweight (Hu et al., 2024): Wu et al. (2024) propose an unbiased watermarking method named Dipmark. Dipmark shuffles all probability masses $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ over the vocabulary set within the probability interval $[0,1]$ based on a key $k$. A hyperparameter $\alpha\in[0,0.5]$ partitions the interval $[0,1]$ into three segments: $[0,\alpha]$, $(\alpha,1-\alpha]$, and $(1-\alpha,1]$. Probabilities in the first segment are set to 0, those in the second remain constant, and those in the third are doubled. Dipmark becomes $\gamma$-reweight when $\alpha=0.5$.
RDW (Kuditipudi et al., 2023): We focus on the RDW method via an inverse transform sampling scheme. Given a uniformly distributed key $k$, RDW first shuffles all probability masses within the interval $[0,1]$, then it randomly samples a value $u\sim\text{U}(0,1)$. Here, $u$ is viewed as the cumulative distribution function value of $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ with respect to the permutation; it is subsequently inverse transformed to generate a token. The probability of the sampled token is reweighted to 1, while the probabilities of all other tokens are reweighted to 0.
3 A Simple Protocol for a Low-entropy Scenario
The low-entropy text refers to a relatively deterministic sequence in natural language. The entropy measures the uncertainty of the probability distribution $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ at a single generation step among the token set $\mathcal{V}$, where low entropy means low uncertainty. For example, in code writing, the structure of a code sequence is regularized where few changes can be made (Lee et al., 2023). More explicitly, for a typical English pangram such as ‘The quick brown fox jumps over the lazy dog’ (Kirchenbauer et al., 2023a), both humans and machines should generate similar if not identical output. For example, when provided with the prompt ‘The quick brown fox jumps over the lazy’, the trained LLaMA-2-7B (Touvron et al., 2023) outputs an empirical probability above 0.8 for the next word ‘dog’.
Problem modeling. Low-entropy scenarios exist in text generation tasks of LLMs. We aim to model a simple problem protocol for the low-entropy generation scenario. For simplicity, we consider the low-entropy scenario where only one token probability is significantly large. Specifically, denote $p_{max}$ as the largest probability of a token in the probability distribution $P_{M}(\cdot|x^{-N_{p}:(t-1)})$. We make an intuitive assumption that except $p_{max}$, other $|\mathcal{V}|-1$ probabilities are small enough to uniformly fill in the remaining $1-p_{max}$ probability value.
@@ 4 Method: Sampling Then Accepting
In this section, we propose the Sampling One Then Accepting (STA-1) method, and discuss detecting the STA-1 generated text using the $z$-test. Theoretically, we show that STA-1 is unbiased and its type II error of the $z$-test has statistical guarantees. Next, we analyze previous unbiased watermarks and STA-1 under the low-entropy protocol in Section 3. We finally introduce Sampling M Then Accepting (STA-M), an extension of STA-1.
4.1 Sampling One Then Accepting
We start by proposing the Sampling One Then Accepting (STA-1) method in Algorithm 1, which is always unbiased and easy to analyze. First, the hash value of the last generated token is computed and employed as the seed of a random number generator (RNG). We use the RNG to divide the token set into a green and a red list (Kirchenbauer et al., 2023a). Next, we sample from the original LLM output distribution (as depicted in Line 4 of Algorithm 1), accept the sampling if the token is in the green list (as depicted in Line 5 and Line 6 in Algorithm 1), sample again if the token is in the red list (as depicted in Line 7 and Line 8 in Algorithm 1), and the second sampling is always accepted.
@@ 4.1 Sampling One Then Accepting
3:     Compute the hash of the last token $x^{t-1}$. Partition the token set $\mathcal{V}$ to form the green $G$ and red $R$ list based on key $k$, the hash, and the proportion $\gamma$
@@
8:        Deny the sampling, sample $x^{t}$ from the distribution $p^{t}$
@@
STA-1 is a simple but effective watermarking method. The properties of STA-1 include: (1) STA-1 is an unbiased watermark; (2) The number of green list tokens in STA-1 generated texts has a lower bound on its mean and an upper bound on its variance, which further provides explicit statistical guarantees for the type II error in the STA-1 detection test; (3) STA-1 has a lower risk for low-entropy generation compared to previous work. In deriving theoretical results, we assume that the key $k$ is randomly sampled from a uniform distribution. Therefore, the random partition of green and red lists associated with this key is also uniform (Kirchenbauer et al., 2023a). We start by analyzing the unbiased characteristic of STA-1.
@@
4.1.1 Statistical Test Guarantees of STA-1

### v3

# Watermarking Low-entropy Generation for Large Language Models: An Unbiased and Low-risk Method

@@ Abstract
Recent advancements in large language models (LLMs) have highlighted the risk of misusing them, raising the need for accurate detection of LLM-generated content. In response, a viable solution is to inject imperceptible identifiers into LLMs, known as watermarks. Our research extends the existing watermarking methods by proposing the novel Sampling One Then Accepting (STA-1) method. STA-1 is an unbiased watermark that preserves the original token distribution in expectation and has a lower risk of producing unsatisfactory outputs in low-entropy scenarios compared to existing unbiased watermarks. In watermark detection, STA-1 does not require prompts or a white-box LLM, provides statistical guarantees, demonstrates high efficiency in detection time, and remains robust against various watermarking attacks. Experimental results on low-entropy and high-entropy datasets demonstrate that STA-1 achieves the above properties simultaneously, making it a desirable solution for watermarking LLMs. Implementation codes for this study are available online. [^1] https://github.com/djwei96/STA
Watermarking Low-entropy Generation for Large Language Models: An Unbiased and Low-risk Method
Minjia Mao1thanks: Preprint. Minjia Mao, Dongjun Wei, and Zeyu Chen contribute equally. Corresponding to Minjia Mao.  Dongjun Wei2  Zeyu Chen1  Xiao Fang1  Michael Chau2 1University of Delaware 2The University of Hong Kong {mjmao,chenze,xfang}@udel.edu dongjun@connect.hku.hk, mchau@business.hku.hk
Table 1: Comparison of the Properties of the Proposed Watermark with Properties of Previous Methods.
Watermark Generation Watermark Detection
Method [Uncaptioned image] Unbiased [Uncaptioned image] Low-risk [Uncaptioned image] Guarantee [Uncaptioned image] Black-box [Uncaptioned image] Efficiency [Uncaptioned image] Robustness
Kirchenbauer et al. (2023a)
Lee et al. (2023)
Hu et al. (2024)
Christ et al. (2023)
Kuditipudi et al. (2023)
Wu et al. (2024)
Ours (STA-1)
@@ 1 Introduction
Large language models (LLMs) are large-scale deep learning models that can understand and generate natural languages by learning from a large amount of textual data. As LLMs can generate content more efficiently at a lower cost compared to humans, the risk of LLMs being employed to generate biased, fake, or malicious content is also increasing (Mirsky et al., 2023; Fang et al., 2024; Pan et al., 2023). To reduce the harm caused by LLMs, it is crucial to identify LLM-generated content precisely and efficiently (Kirchenbauer et al., 2023b). A viable solution is to inject watermarks into LLM-generated text. The watermarked text is imperceptible to humans but detectable by certain models (Liu et al., 2023b). This is achieved by controlling the randomness of the token generation process in LLMs (Kirchenbauer et al., 2023a; Lee et al., 2023), with the randomness kept confidential by LLM owners. In this study, we seek a watermark with the following properties during the generation phase, which are crucial for an effective watermark:
Unbiased: The watermark should adjust the probability distribution while maintaining the same expectation as the unwatermarked distribution, making it impossible to discern between watermarked and unwatermarked text.
Low-risk: The watermark should have a low risk of producing unsatisfactory outputs in low-entropy scenarios (e.g., code generation), where high-probability tokens should be sampled even with watermarks.
Furthermore, the watermark should have the following necessary properties during the detection phase:
Black-box: We do not need prompts or a white-box LLM for detection.
Guarantee: We can have a statistical guarantee on the type II error, where the watermark detection fails to identify a watermarked text.
Efficiency: The detection should only require a low time complexity.
Robustness: The watermark is hard to be removed by watermarking attacks.
However, we find that existing watermarking methods cannot satisfy all these properties simultaneously in the generation and detection phases. In response to these challenges, we propose a novel Sampling One Then Accepting (STA-1) method that can simultaneously achieve these properties. We provide an analysis of previous methods in Appendix A and compare them with the proposed STA-1 method in Table 1. Our proposed STA-1 method can be traced back to the original watermarking method (denoted as KGW) (Kirchenbauer et al., 2023a), where the token set is divided into a green and a red list at each generation step. Instead of raising logits in the green list, STA-1 samples a token from the original probability distribution and accepts it if it is in the green list. If the sampled token is in the red list, it resamples another token and accepts it. We theoretically prove that our STA-1 method is an unbiased watermarking method, which is similar to previous unbiased watermarks Hu et al. (2024); Wu et al. (2024).
The STA-1 method also outperforms other unbiased watermarks in low-entropy scenarios with a lower risk of producing unsatisfactory outputs. Specifically, unsatisfactory outputs in low-entropy scenarios represent that under certain watermark keys, the unbiased watermarking method alters the probability distribution too much such that high-probability tokens cannot be sampled at risk. We prove that STA-1 is less risky than previous unbiased watermarks by analyzing the variance of the probability after altering, using a well-adopted risk-return analysis (Sharpe, 1998).
Another benefit of our proposed method is that STA-1 is a natural extension of KGW that inherits its advantages in the detection phase. Specifically, STA-1 counts the number of green list tokens and employs the $z$-test for watermark detection. The $z$-test naturally eliminates the need for prompts and white-box LLMs in detection (which is required in some previous work Hu et al. (2024)) and only requires $O(m)$ time complexity, where $m$ is the number of tokens. Furthermore, we establish the statistical guarantees for the type II error in watermark detection. These guarantees are related to the Gini index of the probability distribution, a common metric in machine learning (Breiman, 2017), compared to the proposed Spike entropy in KGW.
Additionally, we propose STA-M, an extension of STA-1, by setting up a threshold for entropy in generation (Lee et al., 2023; Wang et al., 2023) and sampling more times for high-entropy steps. Although STA-M is not unbiased theoretically, it allows higher watermark strength with small text quality shifts empirically. Based on the experimental results, we also find that our proposed STA-M method has better robustness compared to KGW against various attacks. Our main contributions can be summarized as follows:
@@
1. We propose STA-1, a novel unbiased watermarking method that has a lower risk theoretically compared to other unbiased watermarks. Moreover, we introduce STA-M, an extension of STA-1 that enhances watermark strength with low text quality shifts.
2. We prove that STA-1 has statistical guarantees for the type II error in its detection based on the widely used Gini index. STA-1 also does not require access to white-box LLMs and only requires $O(m)$ time complexity in detection.
@@
3. Experimental results on low-entropy and high-entropy datasets empirically show that STA-1 is unbiased and has a lower risk of unsatisfactory outputs compared to other unbiased watermarks. Meanwhile, STA-M is more robust against different watermarking attacks than existing methods.
@@ 2 Preliminary
Notations. Let $P_{M}$ denote a pretrained LLM and $\mathcal{V}$ is the overall token set. For simplicity, we use Python-style notation for an ordered token sequence, where $x^{-m:n}=(x^{-m},x^{-m+1},\cdots,x^{n})$, $m$ and $n$ are integers. In a typical LLM generation task, an LLM receives a sequence of $N_{p}+1$ tokens $x^{-N_{p}:0}$, known as a prompt, and outputs a sequence of $T$ tokens $x^{1:T}$ step by step. At step $t$, the probability of each token is given by the conditional distribution $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$. The LLM generation follows an autoregressive fashion, where the joint probability of the generated tokens is $P_{M}(x^{1:T}|x^{-N_{p}:0})=\prod_{t=1}^{T}P_{M}(x^{t}|x^{-N_{p}:(t-1)})$.
When applying watermarking methods, the LLM employs a private key $k$ to adjust the conditional distribution from $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ to $P_{M,w}(x^{t}|x^{-N_{p}:(t-1)};k)$, where $P_{M,w}$ indicates a watermarked LLM and the private key $k$ is randomly selected from a key space $K$ according to a known distribution $P_{K}(k)$. An unbiased watermark requires that the expectation of the watermarked distribution equals that of the original distribution (Hu et al., 2024), defined as follows.
@@
$\mathbb{E}_{k\sim P_{K}(k)}\left[P_{M,w}(x^{t}|x^{-N_{p}:(t-1)};k)\right]$ (1)
$=P_{M}(x^{t}|x^{-N_{p}:(t-1)}),$
@@
One of our main contributions is to show that the risk of unsatisfactory outputs in STA-1 is lower. Here, ‘risk’ is specifically defined for unbiased watermarks in low-entropy scenarios. To support our analysis, we introduce a previous biased watermark KGW (as the backbone of our study) (Kirchenbauer et al., 2023a), alongside other unbiased watermarks including Dipmark (Wu et al., 2024), $\gamma$-reweight (Hu et al., 2024), and RDW (Kuditipudi et al., 2023) in Appendix B.
@@ 3 Method: Sampling Then Accepting
In this section, we first propose the Sampling One Then Accepting (STA-1) method and theoretically show that it is unbiased. We then analyze previous unbiased watermarks alongside STA-1 under a low-entropy protocol, showing that STA-1 has a lower risk of producing unsatisfactory outputs. Next, we explore the detection of STA-1-generated text using the $z$-test and provide a statistical guarantee for its type II error based on the Gini index. Finally, we introduce Sampling M Then Accepting (STA-M), an extension of STA-1.
3.1 Sampling One Then Accepting (STA-1)
We start by proposing the STA-1 method in Algorithm 1, which is always unbiased. We first utilize the last generated token from an LLM to compute its hash value and employ this value as the seed of a random number generator (RNG). We then use the RNG to divide the token set into a green and a red list (Kirchenbauer et al., 2023a). Finally, we sample from the original LLM output distribution (as depicted in Line 4 of Algorithm 1). If the token is in the green list (as shown in Lines 5 and 6 of Algorithm 1), we accept the sample. Otherwise, the token must be in the red list (as depicted in Lines 7 and 8 of Algorithm 1), and we sample a token again, always accepting the second sample.
@@ 3.1 Sampling One Then Accepting (STA-1)
3:     Compute the hash of the last token $x^{t-1}$. Partition the token set $\mathcal{V}$ to form the green $G$ and red $R$ lists based on the key $k$, the hash, and the proportion $\gamma$
@@
8:        Deny the sampling (i.e., $x_{c}^{t}\in$ $R$), sample $x^{t}$ from the distribution $p^{t}$
@@
The STA-1 is a simple yet effective method with many great properties. We begin by analyzing the unbiasedness of STA-1. In the following theorem, we assume that the key $k$ is randomly sampled from a uniform distribution. Consequently, the random partition of the green and red lists associated with this key is also uniform (Kirchenbauer et al., 2023a).
@@
3.1.1 Risk in the Low-Entropy Scenario
The STA-1 method outperforms other unbiased watermarks in generating low-entropy texts, demonstrating a lower risk of producing unsatisfactory outputs. Specifically, the low-entropy text refers to a relatively deterministic sequence in natural language. The entropy measures the uncertainty of the probability distribution $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ at a single generation step among the token set $\mathcal{V}$, where low entropy means low uncertainty. For example, in code writing, the structure of a code sequence is regularized where few changes can be made (Lee et al., 2023). More explicitly, for a typical English pangram such as ‘The quick brown fox jumps over the lazy dog’ (Kirchenbauer et al., 2023a), both humans and machines should generate similar if not identical output. For example, when provided with the prompt ‘The quick brown fox jumps over the lazy’, the trained LLaMA-2-7B (Touvron et al., 2023) outputs an empirical probability above 0.8 for the next token ‘dog’. Such low-entropy scenarios are common in text generation tasks of LLMs. In this paper, we aim to model a simple problem protocol for the low-entropy generation scenario.
Low-entropy Protocol. For simplicity, we consider the low-entropy scenario where only one token probability is significantly large. Specifically, denote $p_{max}$ as the largest probability of a token in the probability distribution $P_{M}(\cdot|x^{-N_{p}:(t-1)})$. We make an intuitive assumption that except $p_{max}$, other $|\mathcal{V}|-1$ probabilities are small enough to uniformly fill in the remaining $1-p_{max}$ probability value.
Previous work claims that unbiased watermarks have no impact on text quality by maintaining the same expectation (Hu et al., 2024). However, we challenge this claim in the low-entropy protocol described above. We show that in such a protocol, unbiased watermarks can still affect text quality because of the risk of unsatisfactory outputs. Consider the following example.
Example 1. Assuming that the token set only consists of two tokens $\mathcal{V}=\{A,B\}$, at a typical step, an LLM outputs the probability of generating $A$ ($p_{A}$) and $B$ ($p_{B}$) as $(p_{A},p_{B})=(0.8,0.2)$. Consider the following two unbiased watermarks. $W_{1}$: with a probability of 0.8 always generating $A$ and with a probability of 0.2 always generating $B$; $W_{2}$: with a probability of 0.5, the probability distribution becomes $(p_{A},p_{B})=(0.9,0.1)$ and with the other probability of 0.5, it becomes $(p_{A},p_{B})=(0.7,0.3)$.
In Example 1, one can view the prompt as ‘The quick brown fox jumps over the lazy’, $A$ as the token ‘dog’, and $B$ as all other tokens. It is easy to show that watermarks $W_{1}$ and $W_{2}$ are both unbiased. However, risk-averse people (Pratt, 1978) will prefer watermark $W_{2}$ because $W_{2}$ does not have a possibility that only $B$ is sampled. $B$ represents unsatisfactory outputs in low-entropy scenarios which could significantly harm text quality, and we want the risk of sampling $B$ to be as low as possible. [^2] We refer readers to Appendix D for a conventional example in finance and a better understanding of the analysis via utility theory. At any generation step, let $x_{max}$ denote the token with the maximum probability $p_{max}$. We measure the risk by the variance (Sharpe, 1998) of $p_{max}^{w,k}$ among watermark keys, where $p_{max}^{w,k}$ denotes the altered value of $p_{max}$ with a watermarking method and a key $k$. We show that STA-1 has a lower risk compared to previous unbiased watermarks in the following theorem. To put it plainly, under the same expectation, the variance of the altered probabilities (risk) by STA-1 is lower.
Theorem 2. Assume $11-\alpha\leq p_{max}<$ 1, where $\alpha$ represents the partition hyperparameter used in Dipmark. For the low-entropy protocol above, the STA-1 method has a lower variance in the probability of generating $x_{max}$ compared to other unbiased methods (including Dipmark, $\gamma$-reweight, and RDW) (Hu et al., 2024; Wu et al., 2024; Kuditipudi et al., 2023). Formally,
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]<\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]$ (2)
$=\mathbb{V}_{k\sim P_{K}(k)}^{\gamma\text{-reweight}}\left[p_{max}^{w,k}\right]<\mathbb{V}_{k\sim P_{K}(k)}^{\text{RDW}}\left[p_{max}^{w,k}\right],$
for any $\alpha\in[0,0.5]$ used in Dipmark.
Proof. See Appendix C.2.
3.1.2 Statistical Test Guarantees